# AI-Driven Grading and Moderation for Collaborative Projects in Computer Science Education

### v1

@@ VI Future Directions
To extend functionality and improve adaptability, several research and engineering enhancements are planned. Each enhancement is an individual research topic or implementation effort. There is a lot of work ahead; we highlight several primary directions that need to be addressed in the future.
@@ VI-A IDE Integration
Develop plugins for popular Integrated Development Environments (IDEs), such as Visual Studio Code, IntelliJ, and Eclipse, to track local coding behavior, file ownership, and collaboration patterns.
This feature would enable seamless data collection from developers’ working environments, offering insight into how and when students contribute to a shared project. By tracking keystrokes, file changes, commit frequencies, and time spent per file, these plugins can attribute contributions more accurately and detect patterns such as code ownership, multitasking, and collaboration bugs. It also enhances the Individual Contribution Analyzer (ICA) by providing rich metadata beyond Git history. Significant challenges include preserving student privacy, managing data volume, and ensuring cross-platform compatibility.
@@ VI-B Multimedia Artifact Analysis
Utilize machine learning models to evaluate non-code artifacts, including screenshots, wireframes, presentations, and videos.
Many software engineering and CS capstone projects include deliverables such as UI mockups, slide decks, and demo videos. Traditional grading overlooks these components or evaluates them manually. By applying computer vision, natural language processing, and audio analysis techniques, the system can automatically assess the clarity, completeness, and professionalism of these artifacts. For example, image recognition could assess consistency in UI wireframes; speech analysis could evaluate clarity and fluency in demo videos. This expansion acknowledges the holistic nature of project work, emphasizing the importance of design and communication skills in addition to coding.
@@ VI-C Enhanced Peer Review
Combine subjective peer feedback with objective ICA metrics to produce more balanced and trustworthy individual assessments.
Peer reviews offer a unique perspective into team dynamics and can capture nuances that automated systems miss, such as leadership, initiative, or teamwork quality. However, they can be biased or inconsistent. By integrating peer evaluations with ICA’s behavioral metrics (e.g., code contributions, task tracking, communication patterns), the platform can cross-validate data, flag anomalies, and calibrate scores. Machine learning models may even weigh peer input differently based on the historical reliability of reviewers, thereby fostering a more trustworthy and transparent system.
@@ VI-D Adaptive Grading Models
Apply reinforcement learning or similar techniques to dynamically adjust grading models based on past performance and instructor feedback.
Current grading systems rely on fixed weights (e.g., 60% project quality, 40% individual contribution). Adaptive grading would instead learn from instructor corrections—such as grade adjustments or override patterns—and iteratively improve the scoring logic. Over time, the system can learn how to handle edge cases, identify scoring inconsistencies, and personalize grading for different course contexts. This reduces manual grading effort and improves grading accuracy, especially in large or diverse classes.
@@ VI-E Cross-Disciplinary Expansion
Adapt the platform for use in disciplines beyond computer science, such as business, design, humanities, and engineering.
Many academic fields assign collaborative group projects but lack structured tools to assess them effectively. By abstracting core features (e.g., contribution tracking, artifact analysis, peer review), the platform can support project-based learning in various domains. For example, in design, it may evaluate the evolution of creative assets; in business, it could track contributions to a business plan or pitch deck. This would require customizing assessment models, interfaces, and rubrics to align with discipline-specific expectations and workflows.
@@ VI-F Bias Audits
Implement regular fairness assessments to detect and mitigate grading bias across gender, language, or demographic groups.
AI-powered assessment systems risk encoding or amplifying biases—especially when evaluating subjective artifacts or using peer input. Conducting systematic audits using fairness metrics (e.g., disparate impact, equal opportunity) helps detect whether certain groups are under-represented or over-represented in top scores. This could involve anonymizing submissions during evaluation, monitoring statistical parity, and incorporating fairness-aware ML models. Transparent reporting and instructor tools to review flagged cases would promote ethical and equitable grading.
@@ VI-G LMS & Rubric Integration
Enable the direct export of grades, comments, and rubrics into Learning Management Systems (LMS) such as Canvas, Moodle, Brightspace, and Blackboard.
For broad classroom adoption, the system must integrate with existing instructor workflows. By aligning with LMS APIs, the platform can automate the delivery of final grades, individualized feedback, and rubric-based assessments, saving time and reducing data entry errors. Furthermore, bidirectional integration may allow the system to pull course context, deadlines, and participation data to inform grading. This streamlining is essential for scaling the use of this approach across institutions and ensuring consistency with institutional grading policies.
@@ VII Conclusion
This paper presents a robust, AI-powered grading system that addresses the key challenges in assessing collaborative group projects within computer science education. By leveraging data from code repositories, communication logs, and automated analysis tools, the system enhances fairness, scalability, and pedagogical insight. Initial trials confirm its efficacy, demonstrating strong alignment with instructor evaluations, reduced grading workload, and improved student trust in assessment transparency. Future expansions will focus on handling non-code contributions, multimedia artifacts, and adaptive grading models applicable across multiple disciplines. These enhancements aim to capture a broader spectrum of collaborative work, ensuring that creativity, communication, and teamwork receive fair recognition alongside technical achievement.

### v2

@@ VI Future Directions
To extend functionality and improve adaptability, several research and engineering enhancements are planned. Each enhancement is an individual research topic or implementation effort. There is much work ahead; we highlight several primary directions that require further attention.
@@ VI-A IDE Integration
Develop plugins for popular Integrated Development Environments (IDEs), such as Visual Studio Code, IntelliJ, and Eclipse, to track local coding behavior, file ownership, and collaboration patterns. This feature would enable seamless data collection from developers’ working environments, offering insight into how and when students contribute to a shared project. By tracking keystrokes, file changes, commit frequencies, and time spent per file, these plugins can more accurately attribute contributions and detect patterns such as code ownership, multitasking, and collaboration bugs. It also enhances the Individual Contribution Analyzer (ICA) by providing rich metadata beyond Git history. Significant challenges include preserving student privacy, managing data volume, and ensuring cross-platform compatibility.
@@ VI-B Multimedia Artifact Analysis
Utilize machine learning models to evaluate non-code artifacts, including screenshots, wireframes, presentations, and videos. Many software engineering and CS capstone projects include deliverables such as UI mockups, slide decks, and demo videos. Traditional grading overlooks these components or evaluates them manually. By applying computer vision, natural language processing, and audio analysis techniques, the system can automatically assess the clarity, completeness, and professionalism of these artifacts. For example, image recognition could assess consistency in UI wireframes; speech analysis could evaluate clarity and fluency in demo videos. This expansion acknowledges the holistic nature of project work, emphasizing the importance of design and communication skills in addition to coding.
@@ VI-C Enhanced Peer Review
Combine subjective peer feedback with objective ICA metrics to produce more balanced and trustworthy individual assessments. Peer reviews offer a unique perspective into team dynamics and can capture nuances that automated systems miss, such as leadership, initiative, or teamwork quality. However, they can be biased or inconsistent. By integrating peer evaluations with ICA’s behavioral metrics (e.g., code contributions, task tracking, communication patterns), the platform can cross-validate data, flag anomalies, and calibrate scores. Machine learning models may even weight peer input differently based on reviewers’ historical reliability, thereby fostering a more trustworthy and transparent system.
@@ VI-D Adaptive Grading Models
Apply reinforcement learning or similar techniques to dynamically adjust grading models based on past performance and instructor feedback. Current grading systems rely on fixed weights (e.g., 60% project quality, 40% individual contribution). Adaptive grading would instead learn from instructor corrections—such as grade adjustments or override patterns—and iteratively improve the scoring logic. Over time, the system can learn how to handle edge cases, identify scoring inconsistencies, and personalize grading for different course contexts. This reduces manual grading effort and improves grading accuracy, especially in large or diverse classes.
@@ VI-E Cross-Disciplinary Expansion
Adapt the platform for use in disciplines beyond computer science, such as business, design, humanities, and engineering. Many academic fields assign collaborative group projects but lack structured tools for effective assessment. By abstracting core features (e.g., contribution tracking, artifact analysis, peer review), the platform can support project-based learning in various domains. For example, in design, it may evaluate the evolution of creative assets; in business, it could track contributions to a business plan or pitch deck. This would require customizing assessment models, interfaces, and rubrics to align with discipline-specific expectations and workflows.
@@ VI-F Bias Audits
Implement regular fairness assessments to detect and mitigate grading bias across gender, language, or demographic groups. AI-powered assessment systems risk encoding or amplifying biases—especially when evaluating subjective artifacts or using peer input. Conducting systematic audits using fairness metrics (e.g., disparate impact, equal opportunity) helps detect whether certain groups are under- or over-represented among top scorers. This could involve anonymizing submissions during evaluation, monitoring statistical parity, and incorporating fairness-aware ML models. Transparent reporting and instructor tools to review flagged cases would promote ethical and equitable grading.
@@ VI-G LMS & Rubric Integration
Enable the direct export of grades, comments, and rubrics into Learning Management Systems (LMS) such as Canvas, Moodle, Brightspace, and Blackboard. For broad classroom adoption, the system must integrate with existing instructor workflows. By aligning with LMS APIs, the platform can automate the delivery of final grades, individualized feedback, and rubric-based assessments, thereby saving time and reducing data-entry errors. Furthermore, bidirectional integration may allow the system to pull course context, deadlines, and participation data to inform grading. This streamlining is essential for scaling the use of this approach across institutions and ensuring consistency with institutional grading policies.
@@ VII Conclusion
This paper presents a robust AI-powered grading system that addresses key challenges in assessing collaborative group projects in computer science education. By leveraging data from code repositories, communication logs, and automated analysis tools, the system enhances fairness, scalability, and pedagogical insight. Initial trials confirm its efficacy, demonstrating strong alignment with instructor evaluations, reduced grading workload, and improved student trust in assessment transparency. Future expansions will focus on handling non-code contributions, multimedia artifacts, and adaptive grading models applicable across multiple disciplines. These enhancements aim to capture a broader spectrum of collaborative work, ensuring that creativity, communication, and teamwork receive fair recognition alongside technical achievement.